Solve the system
a) 4x^2 + 3y^2 = 16
5x+ y = 7
Can you please help me out? Thanks so much in advance:)

Answer by rothauserc(4718)

solve second equation for y
y = 7 - 5x
substitute for y in first equation
4x^2 +3(7-5x)^2 = 16
4x^2 +3(49 -70x +25x^2) = 16
4x^2 +147-210x+75x^2 = 16
79x^2 -210x +131 = 0
this factors into
(79x -131)(x - 1) = 0
x = 131/79 or 1
solve for y in first equation
y = 7 - 5(1) = 2
y = 553/79 - 5(131/79) = -102/79
our solution is
x = 1, y = 2
x = 131/79, y = -102/79